OpenStudy (unofficialllyy):

A cyclist rides his bike at a rate of 5 miles per hour. What is this rate in kilometers per hour? How many kilometers will the cyclist travel in 2 hours? In your computations, assume that 1 mile is equal to 1.6 kilometers. Do not round your answers.

OpenStudy (unofficialllyy):

I need to find the rate and the distance traveled in 2 hours.

OpenStudy (unofficialllyy):

@Directrix

Miracrown (miracrown):

First let's write our speed as a fraction. We are starting with miles PER hour so 5 miles per every 1 hour and now we want to convert this to km/hour. so we need to build a fraction with "miles" in the denominator and "km" in the numerator \[\frac{5 \space miles }{ 1 \space hr } \space \space \times \space \frac{ \space km }{ \space miles }\]

OpenStudy (unofficialllyy):

okay

Miracrown (miracrown):

what values can we use for that fraction? how many kilometers are in 1 mile? (hint) it was given in the problem

OpenStudy (unofficialllyy):

1.6

Miracrown (miracrown):

yup so we can plug that in. When we multiply them together, the "mi" on the top and bottom will cancel \[\frac{ 5 \space \times \space 1.6 \space km}{ 1 \space hr }\] So what value do we obtain in km/hr?

OpenStudy (unofficialllyy):

\[\frac{ 8 km }{ 1 hr}\]

Miracrown (miracrown):

Perfect :D

Miracrown (miracrown):

now let's look at your next part... how many km will you travel in 2 hours?

OpenStudy (unofficialllyy):

So the first part would be 8, then I times it by 2 to get the second part?

OpenStudy (unofficialllyy):

So 16?

Miracrown (miracrown):

Yep, you got it! :D

OpenStudy (unofficialllyy):

Yay!! Thanks!

Miracrown (miracrown):

np :)